OpenStudy (anonymous):

If y = x^2 + x, find the value of the differential dy corresponding to a change in x of dx = .1 when x =2 show work please

OpenStudy (amistre64):

considering dx = \(\Delta x\) ... hmmm

OpenStudy (ash2326):

y=x^2+x first let's differentiate w.r.t to x dy/dx=2x+1 now dy=dx(2x+1) given dx=0.1 x=2 dy=0.1(2*2+1) dy=0.1(5) dy=0.5